# Debrup Das Work done during internship at Rakuten India Enterprise Private Limited IIT Kharagpur

### v2

@@ Abstract
Tool-augmented Large Language Models (TALM) are known to enhance the skillset of large language models (LLM), thereby, leading to their improved reasoning abilities across many tasks. While, TALMs have been successfully employed in different question-answering benchmarks, their efficacy on complex mathematical reasoning benchmarks, and the potential complimentary benefits offered by tools for knowledge retrieval and mathematical equation solving, are open research questions. In this work, we present MathSensei, a tool-augmented large language model for mathematical reasoning. Augmented with tools for knowledge retrieval (Bing Web Search), program execution (Python), and symbolic equation solving (Wolfram-Alpha), we study the complimentary benefits of these tools through evaluations on mathematical reasoning datasets. We perform exhaustive ablations on MATH, a popular dataset for evaluating mathematical reasoning on diverse mathematical disciplines. We also conduct experiments involving well-known tool planners to study the impact of tool sequencing on the model performance. MathSensei achieves 13.5% better accuracy over gpt-3.5-turbo with chain-of-thought on the MATH dataset. We further observe that TALMs are not as effective for simpler math word problems (in GSM-8k), and the benefit increases as the complexity and required knowledge increases (progressively over AQuA, MMLU-Math, and higher level complex questions in MATH). The code and data are available at https://github.com/Debrup-61/MathSensei.
@@
TALM Math Discipline Form Search Python WAlpha Plan Tool-Study
@@
MathSensei
Algebra, Precalculus,
Geometry, Probability,
Number Theory & more
Both Both
@@ 1 Introduction
State-of-the-art Large language models (LLMs), including gpt-3.5-turbo, GPT-4, and open-source counterparts like Llama 2 have demonstrated impressive performance across a broad spectrum of NLP tasks Brown et al. (2020); Radford et al. (2019); Chowdhery et al. (2022); OpenAI (2023). However, their consistent failure on established reasoning dimensions, such as mathematical, commonsense, abductive, and multi-hop reasoning Lu et al. (2023b); Cobbe et al. (2021); Huang and Chang (2023) have led the research community to explore various solutions for enhancing their reasoning abilities. This pursuit has given rise to techniques, such as - (1) intelligent prompting variations, such as chain of thought Wei et al. (2022), program of thought Chen et al. (2022), tree of thoughts Yao et al. (2023), and self-refinement Madaan et al. (2023), (2) program-guided solving that generates python code as intermediate steps and offloads execution to a symbolic interpreter Gao et al. (2023), (3) multi-model interaction frameworks, such as Multi-agent Debate Du et al. (2023); Liang et al. (2023) and Round-Table Conference Chen et al. (2023b), 4) tool-augmented LLMs powered by external symbolic tools, APIs, and libraries Schick et al. (2023); Lu et al. (2023a); Paranjape et al. (2023); Yang and Narasimhan (2023); Xie et al. (2023).
@@
In this work, we study the effectiveness of tool-augmented LLMs (TALM) applied to problems involving mathematical reasoning. Recent advancements in TALM frameworks, such as Chameleon Lu et al. (2023a), OlaGPT Xie et al. (2023), ART Paranjape et al. (2023), and SocraticAI Yang and Narasimhan (2023) have explored the effectiveness of incorporating external tools for solving knowledge-intensive reasoning tasks and fundamental mathematical problems (such as, arithmetic and algebra). However, the effectiveness of TALM framework is yet to be validated on mathematical reasoning tasks involving complex computations. In this context, it is imperative to assess the suitability of specific tool combinations across diverse mathematical domains (e.g., PreAlgebra, Calculus, Geometry, Intermediate Algebra, Probability) at varying levels of difficulty. This motivated us to undertake a thorough evaluation of TALM framework in the context of complex mathematical reasoning tasks. We propose and develop MathSensei, a TALM-based framework, comprising a distinct set of tools (also referred to as modules), combined in a sequential fashion. These modules include LLM-based components, such as -  knowledge retriever (KR[Uncaptioned image]). python code generator (PG[Uncaptioned image]), code refiner (CR[Uncaptioned image]), , and solution generator (SG[Uncaptioned image]); and APIs, such as - Bing-Web-Search-API (BS[Uncaptioned image]) and Wolfram-Alpha-API (WA[Uncaptioned image]). As illustrated in Fig. 1, MathSensei adopts the modular architecture from Chameleon Lu et al. (2023a). Through systematic experiments of MathSensei, we aim to discern the effectiveness of each module in addressing specific types of mathematical problems, having varying levels of difficulty.
@@
Our ablations show complimentary abilities of the modules, effect of ordering and combination (such as setting of WA + BS + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image]) surpassing PG + SG ([Uncaptioned image] + [Uncaptioned image])). This further highlights the need for planning strategies. We evaluate two advanced planning techniques within our pipeline, investigating methodologies such as Plan-And-Solve Lu et al. (2023a) and REACT Yao et al. (2022) with MathSensei.
We make following contributions:
@@
2. Through systematic ablations by varying the set and order of modules in our framework, we observe that complex mathematical problems spanning different domains (such as, algebra, calculus, number theory, and probability from the MATH dataset) can be benefited by certain types, combinations, and order of these modules. We observe that the BS[Uncaptioned image] module outperforms the KR[Uncaptioned image]  module for retrieving relevant knowledge for mathematical problems. The setting of WA + BS + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image]) outperforms PG + SG ([Uncaptioned image] + [Uncaptioned image]), demonstrating that program-guided solving techniques Gao et al. (2023); Drori et al. (2022) may not be universally suitable for all mathematical problems. These findings motivate the necessity of exploiting better planning techniques.
Our best configuration of MathSensei, PG + WA + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image]) achieves an impressive performance accuracy of 47.6 % on the MATH dataset, surpassing gpt-3.5-turbo([Uncaptioned image]) with Chain-of-Thought (CoT) prompting by 13.5% Chen et al. (2023a). The same setting shows a performance gain of +11.6% over GPT-4 (with CoT prompting) on Intermediate Algebra problems. For Precalculus, GPT-4 (with CoT prompting) has an accuracy of 26.7%, which gets improved to 28.9% by our WA + PG + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image]) setting. Improvements on AQuA-RAT and MMLU-Math are lower, $2.4\%$ and $3.3\%$ respectively, showing the efficacy decreases as requirement of external knowledge decreases.
@@ 2 Related Work
Prompting Techniques. Large Language Models (LLMs) employing prompting strategies such as Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (POT) Chen et al. (2022) have demonstrated commendable performance on simple mathematical datasets such as GSM-8K Cobbe et al. (2021). However, their efficacy diminishes for datasets requiring complex computations and advanced mathematical knowledge. For instance, on the MATH dataset, GPT-4 with CoT prompting exhibits a notably low accuracy of $42\%$. Several variations of these strategies have been explored to improve accuracy in reasoning tasks. Madaan et al. (2023) proposed self-refine that involves iteratively refining the initial output by utilizing feedback from the same model. Zhou et al. (2023) employs code-based self-verification, by utilizing python code to check simple constraints that the LLM generated output should satisfy and correcting the output if necessary. Similarly, Progressive-Hint-Prompting Zheng et al. (2023) involves multiple turns of interactions, using previously generated answers as hints for subsequent turns. Similar to POT prompting, PAL (Program Aided language models) Gao et al. (2023) adopts a program-guided solving paradigm. It reads natural language problems, generates programs as intermediate reasoning steps, and delegates the solution step to a runtime environment, such as the Python interpreter. Across 13 natural language reasoning tasks within Big-Bench-Hard Suzgun et al. (2022), they observe that program-guided solving consistently outperforms significantly larger models.
@@
Tool-Augmented LLMs. The emerging trend of tool-augmented LLMs has garnered increasing attention within the research community. Large language models, trained on the objective of next-token prediction, excel at generating tokens based on probabilistic patterns in their training data, making them effective in data-intensive tasks. However, their proficiency falls short in capturing nuanced reasoning or token relationships, particularly in domains like mathematics. Consequently, there are instances or specific question types where it would be advantageous for an LLM to leverage support from specialized tools or modules. For instance, consider a question requiring the solution to the roots of a 4th-degree polynomial. The LLM, upon generating a special token followed by a query, can pause its generation and invoke a mathematics knowledge-base like Wolfram Alpha. Wolfram Alpha, in turn, can utilize its API to process the query and return the answer to the LLM, which can then continue its generation. Toolformer Schick et al. (2023) leverages data annotated with such tool calls (using special tokens for tools) and responses to train language models to employ tools as needed in a self-supervised manner. Similarly, the tool-augmented LLM framework Chameleon Lu et al. (2023a) adopts a plug-and-play approach to utilize tools sequentially. In their setup, the sequence of execution of the tools is predetermined based on a target task; the output of each tool is added to the context for subsequent downstream tools in the pipeline. They perform evaluation on multi-modal knowledge-intensive datasets like ScienceQA, TabMWP. Similarly, frameworks such as ART Paranjape et al. (2023) engage in multi-step reasoning, where each step is linked to a tool call. Utilizing search and code tools, ART tackles various tasks across datasets such as MMLU Hendrycks et al. (2021a)and BigBench Srivastava et al. (2023).
@@ 3.2 Modules
$\bullet$  LLM-based Knowledge Retrieval (KR[Uncaptioned image]) - For this module, we design a prompt to extract relevant knowledge from a pre-trained LLM (taking any one from the list of models mentioned in Table 14) in the form of concepts, formulas, mathematical expressions, theorems, definitions, and hints on how to solve a corresponding mathematical question. An example prompt and output is shown in Table 9A).
$\bullet$  Bing Web Search (BS[Uncaptioned image]) - This module queries the Bing-Web-Search-API ([Uncaptioned image]) to extract the most relevant snippets which may contain similar questions and concepts required for solving a mathematical problem. For similar questions search, we directly query the API ([Uncaptioned image]) with a mathematical question. In case of concepts search, we first use an LLM (either gpt-3.5-turbo ([Uncaptioned image]) or text-davinci-003 ([Uncaptioned image])) to generate a query corresponding to the input question, and then call the API ([Uncaptioned image]) to retrieve relevant concepts (refer to Fig. 2 for an example).
Figure 2: Overview of the BSRefer to caption module; We concatenate the similar questions and concepts (which is then used by a downstream module).
$\bullet$  Wolfram Alpha (WA[Uncaptioned image]) - This module (comprising multiple components) calls the Wolfram-Alpha-API ([Uncaptioned image]) using a query in the Wolfram language, retrieving the mathematical information from this knowledge base and utilizing the capabilities of its computation engine. First we employ an LLM to generate contextualized thoughts. Subsequently, based on the generated thought, the next component formulates a Wolfram code language query (referred to as the “Final Query”). On passing this query as input to the Wolfram Alpha API, we get a JSON dictionary object. We extract all the useful information from this dictionary (using an LLM-based extractor) and add it to the context of next module. An overview of the WA[Uncaptioned image] module is presented in Fig. 3.
@@
$\bullet$  Python Generator+Executor (PG[Uncaptioned image]) - We use an LLM that takes as input the current context as a part of a well-structured prompt (shown in Appendix Fig. 4). The LLM is explicitly instructed to use the sympy library for accessing a set of mathematical operations and data structures required. Based on the prompt, the module generates an (executable) Python code, which on execution returns some output(s) or an error message. We handle syntax errors using two setups:
@@
    Without refinement: Here, if generated code produces syntax errors, we omit the output of PG[Uncaptioned image] from the context for next module.
@@
    Code-Refinement (CR[Uncaptioned image]): Here, we feed the error message along with the incorrect program to a code-fixing LLM which then generates a corrected python code and rationales of fixed errors given as “Errors fixed”. We also add the information of common errors from our qualitative analysis in the system prompt to aid the code refinement process. An output for the code refinement setup from the MATH dataset is presented in Appendix Fig.  5.
@@ 4 Experimental Setup
We first introduce the mathematical datasets used in our study (§4.1), followed by the experiments that we perform with various combinations of modules (§4.2). We use gpt-3.5-turbo ([Uncaptioned image]) as the default LLM in LLM-based modules unless mentioned otherwise. This is mainly because it is more accessible and cheaper compared to models like GPT-4. For querying a search-engine, we use Bing-We-Search-API. Please refer to §A.1 for details about online resources that we use.
@@ 4.2 Experiments
We conduct several experiments by meticulous analysis of individual modules in the domain of complex mathematical reasoning, through systematic ablations on the module sequences. For some of our ablations, we use different variants of OpenAI models, such as text-davinci-002 ([Uncaptioned image]) and text-davinci-003 ([Uncaptioned image]) other than the default gpt-3.5-turbo ([Uncaptioned image]). We also employ models from the Llama family, such as Llama-2-7B ([Uncaptioned image]) and Phind-Code-Llama-34B-V2 ([Uncaptioned image]). We use accuracy as our evaluation metric for comparing different settings. Our experiments enquire the following questions:
$\bullet$  What is the impact of adding LLM generated mathematical knowledge relevant to the question [KR[Uncaptioned image] module] before invoking the Solution Generator module [SG[Uncaptioned image] module]? (§5.1)
$\bullet$  How does the Bing Web Search [BS[Uncaptioned image] module] compare against the LLM-based knowledge generation [KR[Uncaptioned image] module] for the task of adding relevant mathematical knowledge and information to the problem solving process? (§5.1, §5.2)
$\bullet$  What is the utility of augmenting mathematical knowledge-bases, such as Wolfram Alpha [WA[Uncaptioned image] module] with LLMs for solving problems across different levels of complexity? How does it compare against the paradigm of program-guided solving? (§5.3)
$\bullet$  What are the benefits of using program guided complex problem solving [PG[Uncaptioned image] module], and impact of LLM-based code refinement [ CR[Uncaptioned image] module] in case of syntactical errors? (§5.4)
@@
$\bullet$  How to plan effective utilization of these modules? How does non-adaptive planning strategies [Plan-And-Solve] compare against dynamic planning strategies such as [REACT] which uses a thought, action, and observation based mechanism. (Appendix §A.1)

### v3

# Debrup Das Work done during internship at Rakuten India Enterprise Private Limited. Debopriyo Banerjee Somak Aditya Ashish Kulkarni

@@ Abstract
Tool-augmented Large Language Models (TALMs) are known to enhance the skillset of large language models (LLMs), thereby, leading to their improved reasoning abilities across many tasks. While, TALMs have been successfully employed in different question-answering benchmarks, their efficacy on complex mathematical reasoning benchmarks, and the potential complementary benefits offered by tools for knowledge retrieval and mathematical equation solving are open research questions. In this work, we present MathSensei, a tool-augmented large language model for mathematical reasoning. We study the complementary benefits of the tools - knowledge retriever (Bing Web Search), program generator + executor (Python), and symbolic equation solver (WolframAlpha-API) through evaluations on mathematical reasoning datasets. We perform exhaustive ablations on MATH, a popular dataset for evaluating mathematical reasoning on diverse mathematical disciplines. We also conduct experiments involving well-known tool planners to study the impact of tool sequencing on the model performance. MathSensei achieves 13.5% better accuracy over gpt-3.5-turbo with Chain-of-Thought on the MATH dataset. We further observe that TALMs are not as effective for simpler math word problems (in GSM-8K), and the benefit increases as the complexity and required knowledge increases (progressively over AQuA, MMLU-Math, and higher level complex questions in MATH). The code and data are available at https://github.com/Debrup-61/MathSensei
11affiliationtext: IIT Kharagpur
debrupdas@iitkgp.ac.in, saditya@cse.iitkgp.ac.in
22affiliationtext: Rakuten Group, Inc. Rakuten Institute of Technology India
{debopriyo.banerjee,ashish.kulkarni}@rakuten.com
https://debrup-61.github.io/MathSensei.github.io/
@@
{NiceTabular}
|>c|c|c|c|c|c|c|c| TALM Math Discipline Form Search Python WAlpha Plan Tool-Study
@@
MathSensei Algebra, Precalculus, Geometry, Probability, Number Theory & more Both Both
@@ 1 Introduction
State-of-the-art Large language models (LLMs), including gpt-3.5-turbo, GPT-4, and open-source counterparts,
@@
such as Llama 2 have demonstrated impressive performance across a broad spectrum of NLP tasks Brown et al. (2020); Radford et al. (2019); Chowdhery et al. (2024); OpenAI (2023). However, their consistent failure on established reasoning dimensions, such as mathematical, commonsense, abductive, and multi-hop reasoning Lu et al. (2023b); Cobbe et al. (2021); Huang and Chang (2023) have led the research community to explore various solutions for enhancing their reasoning abilities. This pursuit has given rise to techniques, such as - (1) intelligent prompting variations, such as chain of thought Wei et al. (2022), program of thought Chen et al. (2023c), tree of thoughts Yao et al. (2023), and self-refinement Madaan et al. (2023), (2) program-guided solving that generates python code as intermediate steps and offloads execution to a symbolic interpreter Gao et al. (2023), (3) multi-model interaction frameworks, such as Multi-agent Debate Du et al. (2023); Liang et al. (2023) and Round-Table Conference Chen et al. (2023b), 4) tool-augmented LLMs powered by external symbolic tools, APIs, and libraries Schick et al. (2023); Lu et al. (2023a); Paranjape et al. (2023); Yang and Narasimhan (2023); Xie et al. (2023).
@@
In this work, we study the effectiveness of tool-augmented LLMs (TALM) applied to problems involving mathematical reasoning. Recent advancements in TALM frameworks, such as Chameleon Lu et al. (2023a), OlaGPT Xie et al. (2023), ART Paranjape et al. (2023), and SocraticAI Yang and Narasimhan (2023) have explored the effectiveness of incorporating external tools for solving knowledge-intensive reasoning tasks and fundamental mathematical problems (such as, arithmetic and algebra). However, the effectiveness of TALM framework is yet to be validated on mathematical reasoning tasks involving complex computations. In this context, it is imperative to assess the suitability of specific tool combinations across diverse mathematical domains (e.g., PreAlgebra, Calculus, Geometry, Intermediate Algebra, Probability) at varying levels of difficulty. This motivated us to undertake a thorough evaluation of TALM framework in the context of complex mathematical reasoning tasks. We propose and develop MathSensei, a TALM-based framework, comprising a distinct set of tools (also referred to as modules), combined in a sequential fashion. These modules include LLM-based components, such as -  knowledge retriever (KR[Uncaptioned image]). python code generator (PG[Uncaptioned image]), code refiner (CR[Uncaptioned image]), , and solution generator (SG[Uncaptioned image]); and APIs, such as - Bing-Web-Search-API (BS[Uncaptioned image]) and WolframAlpha-API (WA[Uncaptioned image]). As illustrated in Fig. 1, MathSensei adopts the modular architecture from Chameleon Lu et al. (2023a). Through systematic experiments of MathSensei, we aim to discern the effectiveness of each module in addressing specific types of mathematical problems, having varying levels of difficulty.
@@
Our extensive ablations (varying the set and order of modules), show that complex mathematical problems, spanning different subdomains can be benefited by specific types, combinations, and order of the modules. This further highlights the need for planning strategies. We evaluate two advanced planning techniques within our pipeline, investigating methodologies such as Plan-And-Solve Lu et al. (2023a) and REACT Yao et al. (2022) with MathSensei.
We make the following contributions:
@@
2. Through systematic ablations by varying the set and order of modules in our framework, we observe that complex mathematical problems spanning different domains (such as, algebra, calculus, number theory, and probability from the MATH dataset) can be benefited by certain types, combinations, and order of these modules. We observe that the BS module outperforms the KR module for retrieving relevant knowledge for mathematical problems. The setting of WA+BS+SG outperforms PG+SG, demonstrating that program-guided solving techniques Gao et al. (2023); Drori et al. (2022) may not be universally suitable for all mathematical problems. These findings motivate the necessity of exploiting better planning techniques.
Our best configuration of MathSensei, PG+WA+SG achieves an impressive performance accuracy of 47.6 % on the MATH dataset, surpassing gpt-3.5-turbo([Uncaptioned image]) with Chain-of-Thought (CoT) prompting by 13.5% Chen et al. (2023a). The same setting shows a performance gain of +11.6% over GPT-4 (with CoT prompting) on Intermediate Algebra problems. For Precalculus, GPT-4 (with CoT prompting) has an accuracy of 26.7%, which gets improved to 28.9% by our WA+PG+SG setting. Improvements on AQuA-RAT and MMLU-Math are lower, $2.4\%$ and $3.3\%$ respectively, showing the efficacy decreases as requirement of external knowledge decreases.
@@ 2 Related Work
Prompting Techniques. Large Language Models (LLMs) employing prompting strategies such as Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (POT) Chen et al. (2023c) have demonstrated commendable performance on simple mathematical datasets such as GSM-8K Cobbe et al. (2021). However, their efficacy diminishes for datasets requiring complex computations and advanced mathematical knowledge. For instance, on the MATH dataset, GPT-4 exhibits a notably low accuracy of around $50\%$. Several variations of these strategies have been explored to improve accuracy in reasoning tasks. Madaan et al. (2023) proposed self-refine that involves iteratively refining the initial output by utilizing feedback from the same model. Zhou et al. (2024) employs code-based self-verification, by utilizing python code to check simple constraints that the LLM generated output should satisfy and correcting the output if necessary. Similarly, Progressive-Hint-Prompting Zheng et al. (2023) involves multiple turns of interactions, using previously generated answers as hints for subsequent turns. Similar to POT prompting, PAL (Program Aided language models) Gao et al. (2023) adopts a program-guided solving paradigm. It reads natural language problems, generates programs as intermediate reasoning steps, and delegates the solution step to a runtime environment, such as a Python interpreter. Across 13 natural language reasoning tasks within Big-Bench-Hard Suzgun et al. (2023), they observe that program-guided solving consistently outperforms significantly larger models.
@@
Tool-Augmented LLMs. The emerging trend of tool-augmented LLMs has garnered increasing attention within the research community. Large language models, trained on the objective of next-token prediction, excel at generating tokens based on probabilistic patterns in their training data, making them effective in data-intensive tasks. However, their proficiency falls short in capturing nuanced reasoning or token relationships, particularly in mathematical domains. Consequently, there are instances or specific question types where it would be advantageous for an LLM to leverage support from specialized tools or modules. For instance, consider a question requiring the solution to the roots of a 4th-degree polynomial. The LLM, upon generating a special token followed by a query, can pause its generation and invoke a mathematics computing platform WolframAlpha. WolframAlpha, in turn, can utilize its API to process the query and return the answer to the LLM, which can then continue its generation. Toolformer Schick et al. (2023) leverages data annotated with such tool calls (using special tokens for tools) and responses to train language models to employ tools as needed in a self-supervised manner. Similarly, the tool-augmented LLM framework Chameleon Lu et al. (2023a) adopts a plug-and-play approach to utilize tools sequentially. In their setup, the sequence of execution of the tools is predetermined based on a target task; the output of each tool is added to the context for subsequent downstream tools in the pipeline. They perform evaluation on multi-modal knowledge-intensive datasets, such as ScienceQA and TabMWP. Similarly, frameworks such as ART Paranjape et al. (2023) engage in multi-step reasoning, where each step is linked to a tool call. Utilizing search and code tools, ART tackles various tasks across datasets such as MMLU Hendrycks et al. (2021a)and BigBench Srivastava et al. (2023).
@@ 3.2 Modules
$\bullet$  LLM-based Knowledge Retrieval (KR) - For this module, we design a prompt to extract relevant knowledge from a pre-trained LLM (taking any one from the list of models mentioned in Table 6) in the form of concepts, formulas, mathematical expressions, theorems, definitions, and hints on how to solve a corresponding mathematical question. An example prompt and output is shown in Table 19 in Appendix.
$\bullet$  Bing Web Search (BS) - This module queries the Bing-Web-Search-API ([Uncaptioned image]) to extract the most relevant snippets which may contain similar questions and concepts required for solving a mathematical problem. For similar questions search, we directly query the API with the mathematical question. In case of concepts search, we first use an LLM (either gpt-3.5-turbo or text-davinci-003) to generate a query corresponding to the input question, and then call the API to retrieve relevant concepts (refer to Fig. 2 for an example).
Figure 2: Overview of the BS module; We concatenate the similar questions and concepts (which is then used by a downstream module).
$\bullet$  WolframAlpha (WA) - This module (comprising multiple components) calls the WolframAlpha-API using a query in the Wolfram language, retrieving the mathematical information from this knowledge base and utilizing the capabilities of its computation engine. First we employ an LLM to generate contextualized thoughts. Subsequently, based on the generated thought, the next component formulates a Wolfram code language query (referred to as the “Final Query”). On passing this query as input to the WolframAlpha-API, we get a JSON dictionary object. We extract all the useful information from this dictionary (using an LLM-based extractor) and add it to the context of next module. An overview of the WA module is presented in Fig. 3.
@@
$\bullet$  Python Generator+Executor (PG) - We use an LLM that takes as input the current context as a part of a well-structured prompt (shown in Fig. 4). The LLM is explicitly instructed to use the Sympy library for accessing a set of mathematical operations and data structures required. Based on the prompt, the module generates an (executable) Python code, which on execution returns some output(s) or an error message. We handle syntax errors using two setups:
@@
    Without refinement: Here, if generated code produces syntax errors, we omit the output of PG from the context for next module.
@@
    Code-Refinement (CR): Here, we feed the error message along with the incorrect program to a code-fixing LLM which then generates a corrected python code and rationales of fixed errors given as “Errors fixed”. We also add the information of common errors from our qualitative analysis in the system prompt to aid the code refinement process. An output for the code refinement setup from the MATH dataset is presented in Fig. 4 (Appendix).
@@ 4 Experimental Setup
We first introduce the mathematical datasets used in our study (§4.1), followed by the experiments that we perform with various combinations of modules (§4.2). We use gpt-3.5-turbo as the default LLM in LLM-based modules unless mentioned otherwise. This is mainly because it is more accessible and cheaper compared to GPT-4. For querying a search-engine, we use Bing-Web-Search-API. Please refer to A in Appendix for details about online resources that we use.
@@ 4.2 Experiments
We conduct several experiments by meticulous analysis of individual modules in the domain of complex mathematical reasoning, through systematic ablations on the module sequences. For some of our ablations, we use different variants of OpenAI models, such as text-davinci-002 and text-davinci-003 other than the default gpt-3.5-turbo. We also employ models from the Llama family, such as Llama-2-7B and Phind-Code-Llama-34B-V2. We use accuracy as our evaluation metric for comparing different settings. Our experiments enquire the following questions:
$\bullet$  What is the impact of adding LLM generated mathematical knowledge relevant to the question [KR module] before invoking the Solution Generator module [SG module]? (§5.1)
$\bullet$  How does Bing Web Search [BS module] compare against the LLM-based knowledge generation [KR module] for the task of adding relevant mathematical knowledge and information to the problem solving process? (§5.1, §5.2)
$\bullet$  What is the utility of augmenting mathematical knowledge-bases, such as WolframAlpha [WA module] with LLMs for solving problems across different levels of complexity? How does it compare against the paradigm of program-guided solving? (§5.3)
$\bullet$  What are the benefits of using program-guided complex problem solving [PG module], and impact of LLM-based code refinement [CR module] in case of syntactical errors? (§5.4)
@@
$\bullet$  How to plan effective utilization of these modules? How does non-adaptive planning strategies [Plan-And-Solve] compare against dynamic planning strategies such as [REACT] which uses a thought, action, and observation based mechanism. (Appendix A)